An aerospace company has submitted bids on two separate federal government defense contracts, A and B. The company estimates that it has a 60% chance of winning contract A and a 30% chance of winning contract B. If P(B|A)=0.4.

a.What is the probability that the company will win at least one contract?
b.What is the probability that the company will win exactly one contract?
c.What is the probability that the company will win no contract?

Answer by robertb(5830)

a. P(A) + P(B) - P(A and B) = P(A) + P(B) - P(B|A)*P(A) = 0.6 + 0.3 - 0.4*0.6 = 0.66.

b. 0.66 - 0.24 = 0.42

c. 1 - P(A or B) = 1 - 0.66 = 0.34.